# Surface groups among cubulated hyperbolic and one-relator groups

### v1

@@ 4.1 Cubulated hyperbolic groups
(with $g_{i}\in G_{0}$) define a cut set in $\mathrm{Wh}_{X}(Y)$. Since $\mathrm{Wh}_{X}(Y)$ has no 0-cut or 1-cuts, $k\geq 2$. Therefore $Y$ is a $k$-cut in $\widetilde{X}$ of width greater than $C$, contradicting the choice of $C$. ∎
@@
Now consider the case where the JSJ decomposition is non-trivial. Every vertex group $V$ is quasiconvex [7, Proposition 1.2], and in particular finitely generated and hyperbolic. If some vertex group $V$ is not itself free, then by Grushko’s theorem it has a one-ended free factor $V^{\prime}$, which provides a one-ended, quasiconvex subgroup, necessarily of infinite index in $\pi_{1}(X)$. On the other hand, if every vertex group is free, then $\pi_{1}(X)$ is either a surface group or contains a one-ended, quasiconvex subgroup by [52, Theorem 3]. ∎
@@ 4.2 One-relator groups
Combining the result for cubulated hyperbolic groups with recent joint work of the author with Louder [30, 31], and also work of Linton [29], we can also answer Question 0.1 for one-relator groups. The key is Puder’s notion of primitivity rank [37].
@@
Item (i) is a result of the classical theorem of Whitehead that, for non-trivial $w$, $G$ is free if and only if $w$ is primitive [32, Proposition 5.10]. It is easy to see that $\pi(w)=1$ if and only if $w$ is a proper power in $F$; a classical theorem of Karrass–Magnus–Solitar asserts that this is true exactly when $G$ has torsion [32, Proposition 5.17]. Item (iii) is a theorem of Louder and the author [30, Theorem 1.5], while item (iv) is a theorem of Linton [29, Theorem 8.2] (using deep work of Wise [56], as well as further work of Louder and the author [31]). ∎
@@
The final piece of the jigsaw is a recent result of Gardam–Kielak–Logan, who handled the two-generator, one-relator case of Question 0.1 [20, Theorem 1.5].
@@ 5 Questions
Strebel’s theorem asserts that every subgroup of infinite index in an $n$-dimensional Poincaré duality group has cohomological dimension at most $n-1$ [48]. Since groups of cohomological dimension one are precisely free groups [49], Theorems A and D can be thought of as converses to Strebel’s theorem for dimension $n=2$, in the contexts of cubulated hyperbolic and one-relator groups. This is the whole story for one-relator groups, which are always virtually of cohomological dimension 2. However, for cubulated hyperbolic groups, it makes sense to ask whether the converses of the higher-dimensional instances of Strebel’s theorem are also true.
Question 5.1.
Let $X$ be a hyperbolic cube complex with $\operatorname{cd}(\pi_{1}(X))=n>2$. If the cohomological dimension of every subgroup of infinite index in $\pi_{1}(X)$ is less than $n$, does it follow that $\pi_{1}(X)$ must be a $PD_{n}$ group?
As with Theorem A, it is also conceivable that the geometric hypotheses of Question 5.1 can be relaxed.
@@
An answer to Question 5.2 in the case of lattices in products of trees would be interesting. The well known flat-closing conjecture would imply a positive answer.
In a similar vein, removing the cubulation hypothesis from Theorem A would be a major breakthrough.
Question 5.3.
@@
Of course, if every one-ended hyperbolic group had a surface subgroup, then Question 5.3 would have an affirmative answer.
The next question addresses the slight mismatch between the conclusions of our main theorems.
Question 5.4.
@@
Question 5.5.
Let $G$ be a one-ended cubulable group. Does $G=\pi_{1}(X)$, where $X$ is essential and no vertex of $X$ has a link with a separating simplex?
Shepherd’s theorem (Theorem 2.30) implies that there is an $X$ without separating vertices in the links, but does not provide information about higher-dimensional simplices. Regardless of the answer to Question 5.5, it is also important to find an algorithmic condition to recognise whether or not $X$ admits 0- or 1-cuts.

### v2

@@ 4.1 Cubulated hyperbolic groups
(with $g_{i}\in G_{0}$) defines a cut set in $\mathrm{Wh}_{X}(Y)$. Since $\mathrm{Wh}_{X}(Y)$ has no 0-cut or 1-cuts, $k\geq 2$. Therefore $Y$ is a $k$-cut in $\widetilde{X}$ of width greater than $C$, contradicting the choice of $C$. ∎
@@
Now consider the case where the JSJ decomposition is non-trivial. Every vertex group $V$ is quasiconvex [9, Proposition 1.2], and in particular finitely generated and hyperbolic. If some vertex group $V$ is not itself free, then by Grushko’s theorem it has a one-ended free factor $V^{\prime}$, which provides a one-ended, quasiconvex subgroup, necessarily of infinite index in $\pi_{1}(X)$. On the other hand, if every vertex group is free, then $\pi_{1}(X)$ is either a surface group or contains a one-ended, quasiconvex subgroup by [65, Theorem 3]. ∎
@@ 4.2 One-relator groups
Combining the result for cubulated hyperbolic groups with recent joint work of the author with Louder [38, 39], and also work of Linton [37] and Gardam–Kielak–Logan [23], we can also answer Question 0.1 for one-relator groups. The key is Puder’s notion of primitivity rank [48].
@@
Item (i) is a result of the classical theorem of Whitehead that, for non-trivial $w$, $G$ is free if and only if $w$ is primitive [40, Proposition 5.10]. It is easy to see that $\pi(w)=1$ if and only if $w$ is a proper power in $F$; a classical theorem of Karrass–Magnus–Solitar asserts that this is true exactly when $G$ has torsion [40, Proposition 5.17]. Item (iii) is a theorem of Louder and the author [38, Theorem 1.5], while item (iv) is a theorem of Linton [37, Theorem 8.2] (using deep work of Wise [69], as well as further work of Louder and the author [39]). ∎
@@
The final piece of the jigsaw is a recent result of Gardam–Kielak–Logan, who handled the two-generator, one-relator case of Question 0.1 [23, Theorem 1.5].
@@
5 Strebel’s theorem
Let $G$ be an $n$-dimensional Poincaré duality group. Strebel proved that any subgroup of infinite index in $G$ must have strictly smaller cohomological dimension [61], paralleling an older theorem of Whitehead about manifolds [64].
Theorem 5.1 (Strebel theorem).
Let $G$ be a $PD_{n}$ group. If $H$ is a subgroup of infinite index in $G$, then $\mathrm{cd}(H)<n$.
By Stallings’ theorem, Theorem A can be thought of as a converse to Strebel’s theorem when $n=1$ and $G$ is hyperbolic and cubulated. It is therefore natural to wonder whether, if $G$ remains hyperbolic and cubulated, the converse is also true for larger $n$ [^6] This was stated as an open question in an earlier version of this paper. The purpose of this section is to note that this converse fails. I am grateful to Genevieve Walsh and Macarena Arenas for explaining this to me.
The key point is that Strebel’s theorem is not optimal. Indeed, the following stronger result is given as an exercise in Brown’s book [12, Exercise 4, p. 203].
Theorem 5.2 (Strong Strebel theorem).
Let $G$ be a group of type $FP$ and cohomological dimension $n$, and suppose $H^{n}(G,\mathbb{Z}G)$ is finitely generated as an abelian group. If $H$ is a subgroup of infinite index in $G$, then $\mathrm{cd}(H)<n$.
With the stronger theorem in hand, we can provide an example of a 3-dimensional group, which is not a $PD_{3}$ group, but such that every subgroup of infinite index is at most 2-dimensional.
Example 5.3.
Let $L$ be a square-free flag triangulation of the 2-dimensional torus (which exists by a lemma of Dranishnikov [20, Proposition 2.1]). Let $W_{L}$ be the corresponding right-angled Coxeter group, and let $G=[W_{L},W_{L}]$ be the commutator subgroup. Then $G$ is a hyperbolic, cubulated group, acting cocompactly on the Davis cube complex $\widetilde{X}$; let $X$ be the quotient cube complex. The link of every vertex of $\widetilde{X}$, and hence also of $X$, is isomorphic to the torus $L$.
The group $G$ has geometric and cohomological dimension 3, but is not a Poincaré duality group. Indeed, the Gromov boundary $\partial_{\infty}G$ is the Pontryagin sphere [47, Remark 4.4(1)], which is 2-dimensional with $\check{H}^{2}(\partial_{\infty}G;\mathbb{Z})\cong\mathbb{Z}$ and $\check{H}^{1}(\partial_{\infty}G;\mathbb{Z})\cong\mathbb{Z}^{\infty}$ By the Bestvina–Mess theorem [8, Corollary 1.3],
$H^{*}(G;\mathbb{Z}G)\cong\check{H}^{*-1}(\partial_{\infty}G;\mathbb{Z})\,,$
so $G$ has cohomological dimension 3, and every subgroup of infinite index has cohomological dimension at most 2 by Theorem 5.2. But $G$ is not $PD_{3}$, because $H^{2}(G,\mathbb{Z}G)\cong\check{H}^{1}(\partial_{\infty}G;\mathbb{Z})\neq 0$.
Thus, Theorem A should actually be thought of as a conditional converse to Theorem 5.2 in dimension 2.
@@ 6 Questions
As explained in §5, Theorem A can be thought of as a converse to the strong Strebel theorem (Theorem 5.2) in dimension 2, for cubulated hyperbolic groups. Therefore, we can ask whether similar converses hold in higher dimensions.
Question 6.1.
Let $X$ be a hyperbolic cube complex with $G=\pi_{1}(X)$ and $\operatorname{cd}(G)=n>2$. If the cohomological dimension of every subgroup of infinite index in $G$ is less than $n$, does it follow that $H^{n}(G,\mathbb{Z}G)$ is finitely generated as an abelian group?
@@
Combined with Theorem A, the well known flat-closing conjecture would imply a positive answer.
In a similar vein, removing the cubulation hypothesis from Theorem A would be a major breakthrough. The following question reiterates [5, Questions 1.7 and 1.11].
Question 6.3 (Gromov, Whyte).
@@
Of course, if every one-ended hyperbolic group had a surface subgroup, then Question 6.3 would have an affirmative answer.
The next question addresses the slight mismatch between the conclusions of Theorems A and D.
Question 6.4.
@@
Question 6.5.
Let $G$ be a one-ended cubulable group. Is $G\cong\pi_{1}(X)$, where $X$ is essential and no vertex of $X$ has a link with a separating simplex?
Shepherd’s theorem (Theorem 2.31) implies that there is an $X$ without separating vertices in the links, but does not provide information about higher-dimensional simplices. Regardless of the answer to Question 6.5, it is also important to find an algorithmic condition to recognise whether or not $X$ admits 0- or 1-cuts.